Chapter 02, Problem 001 During a hard sneeze, your eyes might shut for 0.56 s. If you are driving a car at 75 km/h during such a sneeze, how far does the car move during that time? the tolerance is +/-2%
media%2F8f7%2F8f72c705-5d0a-4068-b865-8b
media%2F237%2F237ae3bf-3e7c-4904-a9c6-ad
Answer #1

1)

Given:

t = 0.56 s

v = 75 km/h = 75/3.6 m/s = 20.83 m/s

So, d = vt

= 20.83*0.56

= 11.67 m